# Omnipredicting Single-Index Models with Multi-Index Models

### v1

@@ 2.1 Notation
Let $0\leq\bar{L}\leq L$, $R>0$, $\beta=\Omega(\frac{1}{LR})$, and $d,n\in\mathbb{N}$ be given. Let $\mathcal{D}$ be a distribution supported on $\mathcal{X}\times[0,1]$, where $\mathcal{X}\subseteq\mathbb{B}(L)\subseteq\mathbb{R}^{d}$, such that [^4] The observation $\bar{L}\leq L$ follows because $\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\mathbf{x}\mathbf{x}^{\top}\rVert_{\textup{op}}\leq\textup{Tr}(\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\mathbf{x}\mathbf{x}^{\top})=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\left\lVert\mathbf{x}\right\rVert^{2}]\leq L^{2}$.
$\left\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\mathbf{x}\mathbf{x}^{\top}\right]\right\rVert_{\textup{op}}\leq\bar{L}^{2}\leq L^{2}.$
Finally, let $\mathcal{W}:=\mathbb{B}(R)\subseteq\mathbb{R}^{d}$, and let
@@
We will define our loss functions of interest in Section 2.2. Note that no assumptions on $\mathcal{D}$ are made, except that it is supported on $\mathcal{X}$ with radius $L$ and has second moment bound $\bar{L}^{2}$. In particular, Model 1 makes no assumptions about the relationship between the label $y$ and the features $\mathbf{x}$.
We make two additional remarks about Model 1. First, the assumption $\beta=\Omega(\frac{1}{LR})$ is essentially without loss of generality, as otherwise the entire range of $\sigma\in\mathcal{S}$ is $2\beta LR=o(1)$. We make this assumption to simplify our final complexity statements. Second, we introduce the additional second moment bound parameter $\bar{L}$ to more tightly characterize the performance of our population-level omnipredictor in Theorem 4 (and its ERM application, Corollary 7). There are settings where $\bar{L}$ can be up to $\sqrt{d}$ times smaller than $L$, e.g., if $\mathcal{D}_{\mathbf{x}}$ is a well-conditioned Gaussian with bounded mean. On the other hand, in settings where the support of $\mathcal{D}_{\mathbf{x}}$ lies on or close to a low-dimensional subspace (e.g., in mixture models), $L$ and $\bar{L}$ may be more comparable.
@@
Let $\beta>0$, $0\leq\bar{L}\leq L$, $R>0$, and $d,n\in\mathbb{N}$ be given. In the realizable single-index model, we are in an instance of Model 1, with the additional assumption that (following notation from Model 1), for an unknown $(\sigma_{\star},\mathbf{w}_{\star})\in\mathcal{S}\times\mathcal{W}$,
@@ 2.2 Omniprediction preliminaries
In an instance of Model 1, let $\alpha\in(0,\frac{\varepsilon}{6\bar{L}LR^{2}})$. If $p$ is an $\frac{\varepsilon}{2}$-omnipredictor for SIMs (Definition 4) where we make the replacement $\mathcal{S}\leftarrow\mathcal{S}_{\alpha,\alpha+(1-2\alpha LR)\beta}$ in the definition, then it is also an $\varepsilon$-omnipredictor for SIMs using the original $\mathcal{S}$ from Model 1.
@@ 2 $\mathbf{w}_{0}\leftarrow\mathbf{0}_{d}$
Let $\varepsilon\in(0,1)$. In an instance of Model 2, the iterates of Algorithm 1, with $\eta=\frac{1}{\beta\bar{L}^{2}}$ and $T\geq\frac{\beta^{2}\bar{L}^{2}R^{2}}{\varepsilon}$, satisfy
@@
$\leq\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})\left\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\mathbf{x}\mathbf{x}^{\top}\right]\right\rVert_{\textup{op}}\leq\bar{L}^{2}\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D}).$
The second line used the Cauchy-Schwarz inequality, and the third used our second moment bound assumption in Model 2 (cf. Model 1). We further can bound the left-hand side of (23):
@@
The first inequality used Corollary 4, and the second used that $\sigma_{\star}$ is $\beta$-Lipschitz and monotone. Now combining (23), (24), and (25), and using $\eta=\frac{1}{\beta\bar{L}^{2}}$, we have for all $0\leq t<T$ that
$\frac{1}{2\beta}\overline{\ell_{\mathsf{sq}}}(\sigma_{t},\mathbf{w}_{t};\mathcal{D})\leq\frac{\beta\bar{L}^{2}}{2}\left(\left\lVert\mathbf{w}_{t}-\mathbf{w}_{\star}\right\rVert_{2}^{2}-\left\lVert\mathbf{w}_{t+1}-\mathbf{w}_{\star}\right\rVert_{2}^{2}\right).$
Telescoping the above display and using that $\mathcal{W}=\mathbb{B}(R)$, one of the adjusted squared losses $\overline{\ell_{\mathsf{sq}}}(\sigma_{t},\mathbf{w}_{t};\mathcal{D})$ must be bounded by $\varepsilon$ after $T\geq\frac{\beta^{2}\bar{L}^{2}R^{2}}{\varepsilon}$ iterations, else we have the contradiction
$\frac{T\varepsilon}{2\beta}\leq\frac{1}{2\beta}\sum_{0\leq t<T}\overline{\ell_{\mathsf{sq}}}(\sigma_{t},\mathbf{w}_{t};\mathcal{D})\leq\frac{\beta\bar{L}^{2}R^{2}}{2}.$
@@
To interpret Proposition 2, observe that it shows that in the realizable case, $\approx\frac{\beta^{2}\bar{L}^{2}R^{2}}{\varepsilon}$ iterations of $\mathsf{Isotron}$ suffice to produce a SIM $(\sigma_{t},\mathbf{w}_{t})$ that achieves squared loss comparable to the ground truth $(\sigma_{\star},\mathbf{w}_{\star})$, up to error $\varepsilon$. We can convert this bound to an omnigap guarantee as follows.
@@
$\frac{1}{\beta}\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})\leq\mathsf{OG}(\sigma,\mathbf{w};\sigma_{\star},\mathbf{w}_{\star})\leq\mathsf{OG}(\sigma,\mathbf{w})\leq 2\bar{L}R\sqrt{\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})}.$
@@
$\leq\left\lVert\nabla_{\mathbf{w}}\ell_{\mathsf{m},\sigma}(\mathbf{w};\mathcal{D})\right\rVert_{2}\left\lVert\mathbf{w}-\mathbf{w}^{\prime}\right\rVert_{2}\leq 2\bar{L}R\sqrt{\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})}.$
@@
In an instance of Model 2, let $\varepsilon\in(0,LR)$. The iterates of Algorithm 1, with $\eta=\frac{1}{\beta\bar{L}^{2}}$ and $T\geq\frac{4\beta^{2}\bar{L}^{4}R^{4}}{\varepsilon^{2}}$, satisfy $\mathsf{OG}(\sigma_{t},\mathbf{w}_{t})\leq\varepsilon$ for some $0\leq t<T$.
Interestingly, Corollary 5 shows that in the realizable case of Model 2, we can learn a proper $\varepsilon$-omnipredictor (i.e., our predictor is a SIM) using $\approx\frac{\beta^{2}\bar{L}^{4}R^{4}}{\varepsilon^{2}}$ iterations of $\mathsf{Isotron}$. This is consistent with the fact that there exists a proper 00-omnipredictor: the ground truth predictor $\mathbf{x}\to\sigma_{\star}(\mathbf{w}_{\star}\cdot\mathbf{x})$.
@@ 1 Input: Distribution $\mathcal{D}$ from Model 2, iteration count $T\in\mathbb{N}$, step size $\eta>0$
In an instance of Model 1, let $\varepsilon\in(0,LR)$. Algorithm 2, with $T\geq\frac{\bar{L}^{2}R^{2}}{\varepsilon^{2}}$ and $\eta=\frac{R}{\bar{L}T^{-1/2}}$, returns $p$, an $\varepsilon$-omnipredictor for SIMs (Definition 4), using the post-processings $\{k_{\sigma}\}_{\sigma\in\mathcal{S}}$.
@@
$\mathsf{OG}(\sigma_{t},\mathbf{w}_{t};\sigma,\mathbf{w})\leq\frac{1}{2\eta}\left(\left\lVert\mathbf{w}_{t}-\mathbf{w}_{\star}\right\rVert_{2}^{2}-\left\lVert\mathbf{w}_{t+1}-\mathbf{w}_{\star}\right\rVert_{2}^{2}\right)+\frac{\eta\bar{L}^{2}}{2}.$ (27)
@@
$\leq\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[(\sigma_{t}(\mathbf{w}_{t}\cdot\mathbf{x})-y)^{2}\right]\left\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\mathbf{x}\mathbf{x}^{\top}\right]\right\rVert_{\textup{op}}\leq\bar{L}^{2},$
@@
$\frac{1}{T}\sum_{0\leq t<T}\mathsf{OG}\left(\sigma_{t},\mathbf{w}_{t};\sigma,\mathbf{w}\right)\leq\frac{R^{2}}{2\eta T}+\frac{\eta\bar{L}^{2}}{2}=\frac{\bar{L}R}{\sqrt{T}}\leq\varepsilon.$ (29)
@@
Theorem 4 is very general: it says that in the agnostic setting of Model 1, we can define a multi-index model with $\approx\frac{\bar{L}^{2}R^{2}}{\varepsilon^{2}}$ “predictor heads” $\{\sigma_{t},\mathbf{w}_{t}\}_{0\leq t<T}$ that serves as an omnipredictor for SIMs.
The main shortcoming of Theorem 4 is that it is improper. In the realizable case, Corollary 5 shows that this can be overcome with an additional $(\beta\bar{L}R)^{2}$ factor in the iteration count.
@@
In an instance of Model 1, let $\varepsilon\in(0,LR)$, and let $\{\sigma_{t}\}_{0\leq t\leq T-1}$ and $\{\mathbf{w}_{t}\}_{0\leq t\leq T}$ be the output of Algorithm 1 with $T\geq\frac{\bar{L}^{2}R^{2}}{\varepsilon^{2}}$ and $\eta=\frac{R}{LT^{-1/2}}$. Then letting $\tau$ be a uniformly random index in the range $10\leq\tau\leq T-$ 1, and setting $(\hat{\sigma},\hat{\mathbf{w}})\leftarrow(\sigma_{\tau},\mathbf{w}_{\tau})$, we have
@@
In an instance of Model 1, suppose that $\mathcal{D}$ is the uniform empirical distribution over $\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]}\subset\mathcal{X}\times\{0,1\}$, and let $\varepsilon\in(0,LR)$. Algorithm 2, with $T\geq\frac{\bar{L}^{2}R^{2}}{\varepsilon^{2}}$ and $\eta=\frac{R}{\bar{L}T^{-1/2}}$, returns $p$, an $\varepsilon$-omnipredictor for SIMs (Definition 4), using the post-processings $\{k_{\sigma}\}_{\sigma\in\mathcal{S}}$.
@@
$O\left(\left(nd+n\log^{2}(n)\right)\cdot\frac{\bar{L}^{2}R^{2}}{\varepsilon^{2}}\right).$
@@ 4.1 Empirical convergence for BIR
Let $0<\alpha\leq\beta$, $0\leq\bar{L}\leq L$, $R>0$, and $d,n\in\mathbb{N}$ be given. In the agnostic bi-Lipschitz single-index model, we are in an instance of Model 1, except that the link function family $\mathcal{S}$ is replaced with the bi-Lipschitz family $\mathcal{S}_{\alpha,\beta}$ defined in (10).
@@
$\leq O\left(\frac{1}{\sqrt{n}}\left(LR\sqrt{\log\frac{1}{\alpha LR}}+\sqrt{\frac{\beta}{\alpha}}LR+\sqrt{d\log\frac{\beta}{\alpha}}LR.\right)\right)$

### v2

@@ 2.1 Notation
Let $L,R>0$, $\beta=\Omega(\frac{1}{LR})$, and $d,n\in\mathbb{N}$ be given. Let $\mathcal{D}$ be a distribution supported on $\mathcal{X}\times[0,1]$, where $\mathcal{X}\subseteq\mathbb{B}(L)\subseteq\mathbb{R}^{d}$. Finally, let $\mathcal{W}:=\mathbb{B}(R)\subseteq\mathbb{R}^{d}$, and let
@@
We will define our loss functions of interest in Section 2.2. Note that no assumptions on $\mathcal{D}$ are made, except that it is supported on $\mathcal{X}$ with radius $L$. In particular, Model 1 makes no assumptions about the relationship between the label $y$ and the features $\mathbf{x}$. Finally, note that the assumption $\beta=\Omega(\frac{1}{LR})$ is essentially without loss of generality, as otherwise the entire range of $\sigma\in\mathcal{S}$ is $2\beta LR=o(1)$. We make this assumption to simplify our final complexity statements.
@@
Let $\beta>0$, $L,R>0$, and $d,n\in\mathbb{N}$ be given. In the realizable single-index model, we are in an instance of Model 1, with the additional assumption that (following notation from Model 1), for an unknown $(\sigma_{\star},\mathbf{w}_{\star})\in\mathcal{S}\times\mathcal{W}$,
@@ 2.2 Omniprediction preliminaries
In an instance of Model 1, let $\alpha\in(0,\frac{\varepsilon}{6L^{2}R^{2}})$. If $p$ is an $\frac{\varepsilon}{2}$-omnipredictor for SIMs (Definition 4) where we make the replacement $\mathcal{S}\leftarrow\mathcal{S}_{\alpha,\alpha+(1-2\alpha LR)\beta}$ in the definition, then it is also an $\varepsilon$-omnipredictor for SIMs using the original $\mathcal{S}$ from Model 1.
@@ 2 $\mathbf{w}_{0}\leftarrow\mathbf{0}_{d}$
Let $\varepsilon\in(0,1)$. In an instance of Model 2, the iterates of Algorithm 1, with $\eta=\frac{1}{\beta L^{2}}$ and $T\geq\frac{\beta^{2}L^{2}R^{2}}{\varepsilon}$, satisfy
@@
$\leq\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})\left\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\mathbf{x}\mathbf{x}^{\top}\right]\right\rVert_{\textup{op}}\leq L^{2}\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D}).$
The second line used the Cauchy-Schwarz inequality, and the third used our boundedness assumption in Model 2 (cf. Model 1). We further can bound the left-hand side of (23):
@@
The first inequality used Corollary 4, and the second used that $\sigma_{\star}$ is $\beta$-Lipschitz and monotone. Now combining (23), (24), and (25), and using $\eta=\frac{1}{\beta L^{2}}$, we have for all $0\leq t<T$ that
$\frac{1}{2\beta}\overline{\ell_{\mathsf{sq}}}(\sigma_{t},\mathbf{w}_{t};\mathcal{D})\leq\frac{\beta L^{2}}{2}\left(\left\lVert\mathbf{w}_{t}-\mathbf{w}_{\star}\right\rVert_{2}^{2}-\left\lVert\mathbf{w}_{t+1}-\mathbf{w}_{\star}\right\rVert_{2}^{2}\right).$
Telescoping the above display and using that $\mathcal{W}=\mathbb{B}(R)$, one of the adjusted squared losses $\overline{\ell_{\mathsf{sq}}}(\sigma_{t},\mathbf{w}_{t};\mathcal{D})$ must be bounded by $\varepsilon$ after $T\geq\frac{\beta^{2}L^{2}R^{2}}{\varepsilon}$ iterations, else we have the contradiction
$\frac{T\varepsilon}{2\beta}\leq\frac{1}{2\beta}\sum_{0\leq t<T}\overline{\ell_{\mathsf{sq}}}(\sigma_{t},\mathbf{w}_{t};\mathcal{D})\leq\frac{\beta L^{2}R^{2}}{2}.$
@@
To interpret Proposition 2, observe that it shows that in the realizable case, $\approx\frac{\beta^{2}L^{2}R^{2}}{\varepsilon}$ iterations of $\mathsf{Isotron}$ suffice to produce a SIM $(\sigma_{t},\mathbf{w}_{t})$ that achieves squared loss comparable to the ground truth $(\sigma_{\star},\mathbf{w}_{\star})$, up to error $\varepsilon$. We can convert this bound to an omnigap guarantee as follows.
@@
$\frac{1}{\beta}\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})\leq\mathsf{OG}(\sigma,\mathbf{w};\sigma_{\star},\mathbf{w}_{\star})\leq\mathsf{OG}(\sigma,\mathbf{w})\leq 2LR\sqrt{\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})}.$
@@
$\leq\left\lVert\nabla_{\mathbf{w}}\ell_{\mathsf{m},\sigma}(\mathbf{w};\mathcal{D})\right\rVert_{2}\left\lVert\mathbf{w}-\mathbf{w}^{\prime}\right\rVert_{2}\leq 2LR\sqrt{\overline{\ell_{\mathsf{sq}}}(\sigma,\mathbf{w};\mathcal{D})}.$
@@
In an instance of Model 2, let $\varepsilon\in(0,LR)$. The iterates of Algorithm 1, with $\eta=\frac{1}{\beta L^{2}}$ and $T\geq\frac{4\beta^{2}L^{4}R^{4}}{\varepsilon^{2}}$, satisfy $\mathsf{OG}(\sigma_{t},\mathbf{w}_{t})\leq\varepsilon$ for some $0\leq t<T$.
Interestingly, Corollary 5 shows that in the realizable case of Model 2, we can learn a proper $\varepsilon$-omnipredictor (i.e., our predictor is a SIM) using $\approx\frac{\beta^{2}L^{4}R^{4}}{\varepsilon^{2}}$ iterations of $\mathsf{Isotron}$. This is consistent with the fact that there exists a proper 00-omnipredictor: the ground truth predictor $\mathbf{x}\to\sigma_{\star}(\mathbf{w}_{\star}\cdot\mathbf{x})$.
@@ 1 Input: Distribution $\mathcal{D}$ from Model 2, iteration count $T\in\mathbb{N}$, step size $\eta>0$
In an instance of Model 1, let $\varepsilon\in(0,LR)$. Algorithm 2, with $T\geq\frac{L^{2}R^{2}}{\varepsilon^{2}}$ and $\eta=\frac{R}{LT^{-1/2}}$, returns $p$, an $\varepsilon$-omnipredictor for SIMs (Definition 4), using the post-processings $\{k_{\sigma}\}_{\sigma\in\mathcal{S}}$.
@@
$\mathsf{OG}(\sigma_{t},\mathbf{w}_{t};\sigma,\mathbf{w})\leq\frac{1}{2\eta}\left(\left\lVert\mathbf{w}_{t}-\mathbf{w}_{\star}\right\rVert_{2}^{2}-\left\lVert\mathbf{w}_{t+1}-\mathbf{w}_{\star}\right\rVert_{2}^{2}\right)+\frac{\eta L^{2}}{2}.$ (27)
@@
$\leq\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[(\sigma_{t}(\mathbf{w}_{t}\cdot\mathbf{x})-y)^{2}\right]\left\lVert\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}\left[\mathbf{x}\mathbf{x}^{\top}\right]\right\rVert_{\textup{op}}\leq L^{2},$
@@
$\frac{1}{T}\sum_{0\leq t<T}\mathsf{OG}\left(\sigma_{t},\mathbf{w}_{t};\sigma,\mathbf{w}\right)\leq\frac{R^{2}}{2\eta T}+\frac{\eta L^{2}}{2}=\frac{LR}{\sqrt{T}}\leq\varepsilon.$ (29)
@@
Theorem 4 is very general: it says that in the agnostic setting of Model 1, we can define a multi-index model with $\approx\frac{L^{2}R^{2}}{\varepsilon^{2}}$ “predictor heads” $\{\sigma_{t},\mathbf{w}_{t}\}_{0\leq t<T}$ that serves as an omnipredictor for SIMs.
The main shortcoming of Theorem 4 is that it is improper. In the realizable case, Corollary 5 shows that this can be overcome with an additional $(\beta LR)^{2}$ factor in the iteration count.
@@
In an instance of Model 1, let $\varepsilon\in(0,LR)$, and let $\{\sigma_{t}\}_{0\leq t\leq T-1}$ and $\{\mathbf{w}_{t}\}_{0\leq t\leq T}$ be the output of Algorithm 1 with $T\geq\frac{L^{2}R^{2}}{\varepsilon^{2}}$ and $\eta=\frac{R}{LT^{-1/2}}$. Then letting $\tau$ be a uniformly random index in the range $10\leq\tau\leq T-$ 1, and setting $(\hat{\sigma},\hat{\mathbf{w}})\leftarrow(\sigma_{\tau},\mathbf{w}_{\tau})$, we have
@@
In an instance of Model 1, suppose that $\mathcal{D}$ is the uniform empirical distribution over $\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]}\subset\mathcal{X}\times\{0,1\}$, and let $\varepsilon\in(0,LR)$. Algorithm 2, with $T\geq\frac{L^{2}R^{2}}{\varepsilon^{2}}$ and $\eta=\frac{R}{LT^{-1/2}}$, returns $p$, an $\varepsilon$-omnipredictor for SIMs (Definition 4), using the post-processings $\{k_{\sigma}\}_{\sigma\in\mathcal{S}}$.
@@
$O\left(\left(nd+n\log^{2}(n)\right)\cdot\frac{L^{2}R^{2}}{\varepsilon^{2}}\right).$
@@ 4.1 Empirical convergence for BIR
Let $0<\alpha\leq\beta$, $L,R>0$, and $d,n\in\mathbb{N}$ be given. In the agnostic bi-Lipschitz single-index model, we are in an instance of Model 1, except that the link function family $\mathcal{S}$ is replaced with the bi-Lipschitz family $\mathcal{S}_{\alpha,\beta}$ defined in (10).
@@
$\leq O\left(\frac{1}{\sqrt{n}}\left(LR\sqrt{\log\frac{1}{\alpha LR}}+\sqrt{\frac{\beta}{\alpha}}LR+\sqrt{d\log\frac{\beta}{\alpha}}LR\right)\right).$